Fermi National Laboratory


Accelerator Update for August 29, 2005 - August 31, 2005

Monday, August 29
The day shift began with Operations monitoring store 4358, the Recycler Ring (RR) stash at 68.6 E10, the Antiproton Source (Pbar) stack at 152.2 E10, and with MiniBooNE, NuMI, and the Meson MCenter taking beam.

Operations put the Recycler into shot setup at 11:48 AM to stash antiprotons.

Pbar returns to stacking at 1:24 PM, beam returns to MCenter.

At 2:44 PM, all NuMI BPMs were reading back zero, this caused NuMI to run at lower intensities until 4:11 PM when the problem was resolved.

Tuesday, August 30
The midnight shift began with Operations monitoring store 4358, the Recycler Ring (RR) stash at 114.6 E10, the Pbar stack at 122.7 E10, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations terminated store 4358 at 8:10 AM.

Shot setup begins at 10:08 AM.

Operations established store 4359 at 11:47 AM with an initial luminosity of 125.4 E10.

Pbar resumed stacking at 11:49 AM.

Operations put the Recycler into shot setup at 3:37 PM to stash antiprotons.

Pbar returns to stacking at 5:24 PM, beam returns to MCenter.

Operations put the Recycler into shot setup at 11:02 PM to stash antiprotons.

Wednesday, August 31
The midnight shift began with Operations monitoring store 4359, the Recycler Ring (RR) stash at 142.1 E10, and with MiniBooNE, NuMI, and MCenter taking beam.

Pbar returns to stacking at 12:41 AM, beam to MCenter.

Operations put the Recycler into shot setup at 5:53 AM to stash antiprotons.

Recycler transfers were completed at 6:30 AM, we have a new record for Pbars in the recycler (191.0 E10).

Pbar returns to stacking at 6:35 AM, beam to MCenter.

Pbar turned off at 7:00 AM to prepare for access.

Booster, MiniBooNE, NuMI, and MCenter turned off at 7:30 AM to prepare for access.

The Plan for Wednesday
The plan for today is to allow access to Pbar, Booster, MiniBooNE, NuMI, and switchyard. No access to Tevatron and Main Injector. Current store will be terminated at 7:30 PM, followed by CDF and D0 accesses. Shot setup will occur on the Thursday midnight shift.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/31/2005   email Fermilab